# Graded Projection Recursion (GPR): A Framework for Controlling Bit-Complexity of Algebraic Packing

### v3

@@ 22 Discussion and conclusions
The framework is strongest when (1) the host algorithm is GEMM-dominant in a blocked/recursive form, and (2) a projection discipline exists that keeps the relevant bands uniformly bounded. The main improvements ahead are sharpening constants, enlarging feasible regimes under more structure assumptions, and exploring alternative extractor mechanisms with better headroom tradeoffs.
@@ Appendix B Exact dyadic-lattice extractor specification (integral mode)
This appendix gives a concrete, implementable specification of the dyadic-lattice rounding operator $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ and the sieve extractor $\mathcal{S}_{\beta}(\cdot)$ used in Part I. The purpose is to eliminate any ambiguity about “exactness” in the integral semantics: all operations below are exact on the declared success set (master gap holds with slack), and any failure of the gap condition can be detected.
@@ B.2 Exact nearest-integer rounding of dyadic rationals
We use a deterministic nearest-integer rounding $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ with a fixed tie-breaking rule. Because the paper’s extraction guarantees use strict slack, tie-breaking is irrelevant on the success set, but we specify it anyway for completeness (ties-to-even is acceptable).
Definition B.2 (Exact rounding of an integer divided by a power of two).
Let $N\in\mathbb{Z}$ and $s\in\mathbb{Z}_{\geq 0}$. Define $\mathrm{RoundPow2}(N,s):=\left\lfloor N/2^{s}\right\rceil$ computed with $N=\mathrm{sgn}\cdot A$, where $\mathrm{sgn}\in\{+1,-1\}$ and $A=\lvert N\rvert\in\mathbb{Z}_{\geq 0}$:
@@
so $0\leq r<2^{s}$. Let $h:=2^{s-1}$ for $s\geq 1$ (and $h:=0$ for $s=0$). Define:
$\tilde{q}~:=~\begin{cases}q,&r<h,\\ q+1,&r>h,\\ q~\text{or}~q+1~\text{(tie-breaking)},&r=h.\end{cases}$ (256)
@@
In code, this is implemented by bit shifts and bit masks on the absolute value (using a wider fixed-width type such as __int128 as needed), with a deterministic rule for the $r=h$ tie case.
@@ B.3 Exact coefficient rounding $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$
B.4 Exact sieve extractor $\mathcal{S}_{\beta}(\cdot)$
@@
All steps are exact in $\mathcal{L}_{D}$ (i.e., dividing by $\beta=2^{b}$ is an exponent-offset update) and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma. Therefore the implemented $\mathcal{S}_{\beta}$ satisfies the mathematical sieve definition. ∎
@@ B.5 Certified base choice and failure policy
and we do not treat $\delta$ as a tuning parameter. At every recursion node we require the sufficient-base inequality from Theorem  7.1 with $\delta=\delta_{0}$:
@@
When Eq. 263 holds and the normalization bounds $\|A^{\prime}_{i,:}\|_{2}\leq 1$, $\|B^{\prime}_{:,j}\|_{2}\leq 1$ are satisfied, extraction is exact because the master gap in Lemma  5.3 is open by at least $\delta_{0}$.
@@
$g_{\min}$ $:=$ $\min_{r:a_{r}<\infty,~b_{r}<\infty}(a_{r}+b_{r}),$ (266)
@@
with the convention that if the set $\{r:a_{r}<\infty,~b_{r}<\infty\}$ is empty (i.e., every dot-product term is zero) then $\sigma:=0$. This ensures that every dot-product term $A^{\prime}_{ir}B^{\prime}_{rj}$ lies in the scaled coefficient lattice:
@@
If the node certificate fails (either the normalization bounds are violated, or Eq. 271 fails even after the one-step upgrade), then the implementation must not apply the sieve at that node. Instead it must treat this as leaving the declared exact-execution regime (e.g., due to overflow, truncation below lattice resolution, or a bug) and must fall back to a conventional exact multiplication on the original subproblem, i.e., call ClassicalMM(A,B)(A,B) (Algorithm  3) and return its result. (An optimized implementation may instead increase the staged-precision budget / word width and restart the node; see Algorithm  2 and Theorem  9.2.) With this policy, the algorithm never returns an incorrect extracted value, i.e., failure affects only performance, not correctness.
@@
Input: $A,B\in\mathcal{L}_{D}^{m\times m}$
Output: $C=AB$ exactly
@@
       $s\leftarrow s+A_{ir}B_{rj}$
    $C_{ij}\leftarrow s$
return $C$
Algorithm 3 ClassicalMM(A,B)(A,B) (exact dyadic fallback)

### v4

@@ 22 Discussion and conclusions
The framework is strongest when (1) the host algorithm is GEMM-dominant in a blocked/recursive form, and (2) a projection discipline exists that keeps the relevant bands uniformly bounded. Although the algorithmic structure presented here is sufficient to achieve the principal objective of rigorous and honest near-quadratic scaling, there is need to explore refactorizations of the algorithm to maximize practical performance. This includes relatively obvious space and computation reuse among the four recursion, and extends to restructured “foldings” to optimize the mapping of computations to hardware (serial and/or parallel).
@@ Appendix B Exact dyadic-lattice extractor specification (integral mode)
This appendix gives a concrete, implementable specification of the dyadic-lattice rounding operator $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ and the scaled sieve extractor $\mathcal{S}_{\beta,\sigma}(\cdot)$ used in Part I. The purpose is to eliminate any ambiguity about “exactness” in the integral semantics: all operations below are exact on the declared success set (master gap holds with slack), and any failure of the gap condition can be detected.
@@ B.2 Exact nearest-integer rounding of dyadic rationals
We use a deterministic nearest-integer rounding $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ with ties-to-even tie-breaking. Because the paper’s extraction guarantees use strict slack, tie-breaking is irrelevant on the success set, but we specify it anyway for completeness.
Definition B.2 (Exact rounding of an integer divided by a power of two (ties-to-even)).
Let $N\in\mathbb{Z}$ and $s\in\mathbb{Z}_{\geq 0}$. Define $\mathrm{RoundPow2}(N,s):=\left\lfloor N/2^{s}\right\rceil$ using ties-to-even. Write $N=\mathrm{sgn}\cdot A$, where $\mathrm{sgn}\in\{+1,-1\}$ and $A=\lvert N\rvert\in\mathbb{Z}_{\geq 0}$. If $s=0$, return $\mathrm{RoundPow2}(N,0):=N$. Otherwise ($s\geq 1$), define:
@@
so $0\leq r<2^{s}$, and let $h:=2^{s-1}$. Define:
$\tilde{q}~:=~\begin{cases}q,&r<h,\\ q+1,&r>h,\\ q,&r=h~~\text{and}~~q\equiv 0~(\mathrm{mod}~2),\\ q+1,&r=h~~\text{and}~~q\equiv 1~(\mathrm{mod}~2).\end{cases}$ (279)
@@
In code, this is implemented by bit shifts and bit masks on the absolute value (using a wider fixed-width type such as __int128 as needed), with a deterministic ties-to-even rule for the $r=h$ tie case.
@@ B.3 Exact coefficient rounding $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$
B.4 Exact scaled sieve extractor $\mathcal{S}_{\beta,\sigma}(\cdot)$
@@
All steps are exact in $\mathcal{L}_{D}$ (i.e., dividing by $\beta=2^{b}$ is an exponent-offset update) and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma. Therefore the implemented $\mathcal{S}_{\beta,\sigma}$ satisfies the mathematical sieve definition. ∎
@@ B.5 Certified base choice and failure policy
and we do not treat $\delta$ as a tuning parameter. At every recursion node we require the sufficient-base inequality from Theorem 7.1 with $\delta=\delta_{0}$:
@@
When Eq. 286 holds and the normalization bounds $\|A^{\prime}_{i,:}\|_{2}\leq 1$, $\|B^{\prime}_{:,j}\|_{2}\leq 1$ are satisfied, extraction is exact because the master gap in Lemma 5.3 is open by at least $\delta_{0}$.
@@
$g_{T}$ $:=$ $\min_{r:a_{r}<\infty,~b_{r}<\infty}(a_{r}+b_{r}),$ (289)
$g_{U}$ $:=$ $\min_{r\in\{1,\dots,m/2\}:a_{r}<\infty,~b_{r+m/2}<\infty}(a_{r}+b_{r+m/2}),$ (290)
$g_{\min}$ $:=$ $\min\{g_{T},~g_{U}\},$ (291)
@@
with the conventions that if an index set is empty we interpret the corresponding minimum as $\infty$, and if both minima are $\infty$ (i.e., all dot-product terms contributing to the middle band $T$ and the upper band $U$ vanish) then we set $\sigma:=0$. This ensures that every dot-product term contributing to $T$ and $U$ under the local packing (i.e., $A^{\prime}_{ir}B^{\prime}_{rj}$ and $A^{\prime}_{ir}B^{\prime}_{r+m/2,j}$) lies in the scaled coefficient lattice:
@@
If the node certificate fails (either the normalization bounds are violated, or Eq. 296 fails even after the one-step upgrade), then the implementation must not apply the sieve at that node. Instead it must treat this as leaving the declared exact-execution regime (e.g., due to overflow, truncation below lattice resolution, or a bug) and must fall back to a conventional exact multiplication on the original subproblem, i.e., call ClassicalMM(A,B)(A,B) (Algorithm 4) and return its result. (An optimized implementation may instead increase the staged-precision budget / word width and restart the node; see Algorithm 2 and Theorem 9.2.) With this policy, the algorithm never returns an incorrect extracted value, i.e., failure affects only performance, not correctness.
@@
Algorithm 4 ClassicalMM(A,B)(A,B) (exact dyadic fallback)
$A,B\in\mathcal{L}_{D}^{m\times m}$.
$C=AB$ exactly.
@@
     $s\leftarrow s+A_{ir}B_{rj}$.
   end for
   $C_{ij}\leftarrow s$.
  end for
end for
return $C$.